OpenStudy (anonymous):

Solve algebraically for x: x + 2/6 = 3/x-1

OpenStudy (anonymous):

All I know is the (x+2)(x-1)=0 part then x+2=0 negative 2 on both sides and x-1=0 positive one on both sides

OpenStudy (anonymous):

tell me - shoulld it be (x+2) / 6 = 3/(x-2) ?

OpenStudy (anonymous):

sorry for the confusion. ill just attach an image to show you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i assumed it was x + 6/2 thats why a took so long first we need to multiply thru by the lcm which is 6(x-2): (x+2(x-2) = 18 x^2 - 4 = 18 x^2 = 22 so x = sqrt22 or -sqrt22

OpenStudy (anonymous):

aww - sorry i made a mistake its (x+2)(x-1) = 18 x^2 + x -2 = 18 x^2 + x - 20 = 0 (x+5)(x-4) = 0 x = 4 or -5

OpenStudy (anonymous):

hmm its alright, thank you very much!

OpenStudy (anonymous):

ok - this answer checks out if you plug in 4 and -5 into the original equation plz click on good answer to close this question - unless u have any more queries of course